OpenStudy (anonymous):

The difference of two numbers, a and b, is 21. The difference of 5 times a and two times b is 18. What are the values of a and b?

OpenStudy (anonymous):

18 ?= 18

OpenStudy (anonymous):

@Kemist

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Set up two equations: a - b = 21 [The difference of two numbers, a and b, is 21] 5a - 2b = 18 [The difference of 5 times a and two times b is 18] Use the first equation to express a in terms of b by adding b to both sides: a = b + 21 Substitute (b + 21) for a in the second equation: 5(b + 21) - 2b = 18 Multiply out: 5b + 105 - 2b = 18 Simplify: 3b + 105 = 18 Subtract 105 from both sides: 3b = -87 Divide both sides by 3: b = -29 Substitute -29 for b in the first equation to solve for a: a - (-29) = 21 Simplify: a + 29 = 21 Subtract 29 from both sides: a = -8 To check, substitute the values -8 and -29 for a and b in the second equation: 5(-8) - 2(-29) ?= 18 Multiply out: -40 + 58 ?= 18 Simplify: 18 = 18